OpenStudy (anonymous):

Double integral of dxdy/(3+x-y)^2

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\int\limits_{0}^{x+1} \frac{ dxdy }{ (3+x-y)^2 }\]

OpenStudy (amistre64):

let (3-y) be some constant value wrt.x

OpenStudy (amistre64):

\[\int \frac{1}{(k+x)^2}~dx\]

OpenStudy (anonymous):

∫1(k+x)2 dx

OpenStudy (amistre64):

you sure your limits are correct? dxdy should amount to a limit of x from a to f(y)

OpenStudy (anonymous):

they gave me the domain

OpenStudy (anonymous):

what i don't understand is this : \[\int\limits_{0}^{x+1} \frac{ dy }{ (3+x-y)^2 }= \frac{ 1 }{ 3+x-y } from 0 \to x+1\]

OpenStudy (anonymous):

I think you are correct

OpenStudy (anonymous):

y from 0 to x+1

OpenStudy (anonymous):

Here 3+x assumed to be constant

OpenStudy (anonymous):

is there a formula for this? i don't have it in my book..

OpenStudy (amistre64):

they gave you the domain, and you deduced the limits? and came up with\[\Large\int\limits_{1}^{2}\int\limits_{0}^{\color{red}{x+1}} \frac{ 1}{ (3+x-y)^2 }~\color{red}{dx}~dy\]

OpenStudy (anonymous):

yes D: x=1 x=2, y=x+1

OpenStudy (amistre64):

if we solve this as is ... then for dy we have all x parts and its just y

OpenStudy (anonymous):

the integrals are reversed

OpenStudy (anonymous):

first dy and after dx...

OpenStudy (amistre64):